# Covariate-Adaptive Randomization in Clinical Trials without Inflated Variances

### v2

@@ 2.2 Theoretical Properties
If $\textsf{E}Z^{2}<\infty$, then
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=$ $\sum_{i=1}^{n}\big((T_{i}-1)\widetilde{Z}_{i}-\textsf{E}[(\ell_{i}-\rho)\widetilde{Z}_{i}|\bm{\Lambda}_{i-1}]\big)$
$+O(n^{\gamma/2}+n^{1-\gamma})\text{ in }L_{r}$ (2.10)
for all $r>0$, where $\widetilde{Z}_{i}=Z_{i}-\textsf{Pj}[Z_{i}|\bm{\phi}(\bm{X}_{i})]$.
If $\textsf{E}Z^{2}<\infty$, $\textsf{E}W^{2}<\infty$, $\textsf{E}W=0$ and $1/2<\gamma<1$, then (2.8) holds,
and
$\textsf{E}\Big(\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\Big)^{2}\leq C_{\phi}\textsf{E}Z^{2}\text{ with }C_{\phi}\text{ does not depend on }Z.$ (2.12)
Remark 2.2.
If $\ell(x)$ is thrice differentiable at $x=0$, and $\ell^{\prime\prime}(0)=0$ or $\textsf{E}[(\bm{\phi}(\bm{X}))^{\prime}\bm{\phi}(\bm{X})\widetilde{Z}]=\bm{0}$, then $n^{1-\gamma}$ in ((ii)) can be repalced by $n^{1-3\gamma/2}$, and (2.8), (2.11), (2.12) remain true when $1/3<\gamma<1$.
Finally, we propose an allocation function such that (2.8) and (2.11) hold for all $0<\gamma<1$. The function is defined by
where $\lambda>0$ is a tuning parameter. In (2.13), $\Phi(-x+u_{\rho/2})$, $\Phi(-x+u_{(\rho+1)/2})$ can be replaced by $\alpha\rho$, $1-(1-\alpha)\rho$, respectively, $0<\alpha<1$.
Theorem 2.3.
Suppose Assumption 2.1, and that 2.2 is satisfied for all $r_{0}>2$, and $\textsf{E}Z^{2}<\infty$. Let $\ell(x)$ be defined as in (2.13). Then
@@
for all $r>0$. Moreover, (2.8), (2.11) and (2.12) hold for all the case of $0<\gamma<1$ if $\textsf{E}W^{2}<\infty$ and $\textsf{E}W=0$.
@@
Let $D_{n}=\sum_{i=1}^{n}(T_{i}-\rho)$, $D_{n}(l,k_{1}^{\ast})=(T_{i}-\rho)I\{X_{i,l}=x_{l}^{k_{l}^{\ast}}\}$, $D_{n}(k_{1}^{\ast},\ldots,k_{p}^{\ast})=(T_{i}-\rho)I\{X_{i,1}=x_{1}^{k_{1}^{\ast}},\ldots,X_{i,p}=x_{p}^{k_{p}^{\ast}}\}$. Then
@@
Thus, when $w_{s}\neq 0$, $D_{n}^{2}(k_{1},\ldots,k_{p})=O_{P}(n^{\gamma/2})$ and so for any $m(\bm{X})$, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})=O_{P}(n^{\gamma/2})$. When $w_{s}=0$, we will have
$D_{n}(l,k_{l})=$ $O_{P}(n^{\gamma/2}),\text{ if }w_{m,l}\neq 0,$
@@
with $\vec{\sigma}_{m}^{2}=\vec{\sigma}_{m-\textsf{E}m}^{2}+\vec{\sigma}_{\textsf{E}m}^{2}=\vec{\sigma}_{m-\textsf{E}m}^{2}\leq\rho(1-\rho)\textsf{Var}(m(\bm{X}))$ since $1\in Span(\bm{\phi})$. In particular,
@@
When $\rho=1/2$, the allocation probability is similar to that of Zhang, (2021), where (2.16) is obtained, but the properties (2.17) and (2.18) are not founded.
When $\rho=1/2$ and $\gamma=0$, our CRA procedure is similar to that of Hu and Hu, (2012). The theory of this kind of CAR procedures can be found in Hu and Zhang, (2020) and Zhang, (2023). It is shown that (2.17) still holds, but $\vec{\sigma}_{m}^{2}$ has no closed from and may exceed $\textsf{E}[m^{2}(\bm{X})]/4$ unless $\textsf{E}[m(\bm{X})|\bm{\phi}(\bm{X})]\in Span\{\bm{\phi}(\bm{X})\}$. In particular, when $w_{s}=w_{0}=0$, the procedure is Pocock and Simon, (1975)’s marginal procedure. Zhao et al., (2024) demonstrated via examples that along some directions, there is an increased variance of within-stratum imbalances under Pocock and Simon, (1975)’s procedure compared to the simple randomization.
@@
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{2}]=O\big(n^{\frac{1}{r_{0}-1}}\big)$, and $O\big(n^{\frac{8}{(r_{0}+1)^{2}}}\big)$ when $r_{0}\geq 3$.
@@
If $\textsf{E}Z^{2}<\infty$, $\textsf{E}W^{2}<\infty$, $\textsf{E}W=0$, $\textsf{E}[Z|\bm{\phi}(\bm{X})]\in Span\{\bm{\phi}(\bm{X})\}$, and $\big(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})]\big)^{2}\in Span\{\bm{\phi}(\bm{X})\}$ when $\rho\neq 1/2$, then
@@ 3 An application
As an application, we study hypothesis testing of treatment effects under the proposed covariate-adaptive procedure. For each unit $i$, denote $Y_{i}(1)$ and $Y_{i}(2)$ as the potential outcomes under the treatments $1$ and $2$, respectively, with $\textsf{E}[Y_{i}(t)]=\mu_{t}$ and $\textsf{Var}(Y_{i}(t))=\sigma_{t}^{2}$, $t=1,2$. The observed outcome is $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(2).$ The treatment effect is defined as $\tau=\textsf{E}\{Y_{i}(1)\}-\textsf{E}\{Y_{i}(0)\}$. Also, denote $\bm{X}_{i}=(X_{i,1},\ldots,X_{i,p})$ as the covariate of unit $i$. Assume that $(Y_{i}(0),Y_{i}(1),\bm{X}_{i})$, $i=1,\ldots,n$, are independent and identically distributed as $(Y(0),Y(1),\bm{X})$. We suppose that the units are randomized sequentially by the proposed CAR procedure with feature $\bm{\phi}(\bm{X})$. We consider the test of
@@
Since $\overline{Y}_{n,1}=\frac{1}{N_{n,1}}\sum_{i=1}^{n}T_{i}Y_{i}$, $\overline{Y}_{n,2}=\frac{1}{N_{n,2}}\sum_{i=1}^{n}(1-T_{i})Y_{i}$, $\widehat{\sigma}_{n,1}^{2}=\frac{1}{N_{n,1}-1}\sum_{i=1}^{n}T_{i}(Y_{i}-\overline{Y}_{n,1})^{2}$, $\widehat{\sigma}_{n,2}^{2}=\frac{1}{N_{n,2}-1}\sum_{i=1}^{n}(1-T_{i})(Y_{i}-\overline{Y}_{n,2})$ are consistent estimators of $\mu_{1}$, $\mu_{2}$, $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$, respectively, where $N_{n,1}=\sum_{i=1}^{n}T_{i}$ and $N_{n,2}=n-N_{n,1}$, we consider the classical test statistic
@@
Suppose that the proposed CAR procedure with a feature $\bm{\phi}(\bm{X})$ is applied to balance covariates with Assumptions 2.1, 2.2 for all $r_{0}>2$ being satisfied. Then
@@
$\sigma_{\tau}^{2}=\frac{\sigma_{1}^{2}}{\rho}+\frac{\sigma_{2}^{2}}{1-\rho}-\rho(1-\rho)\sigma_{Pj}^{2}=\frac{\widetilde{\sigma}_{1}^{2}}{\rho}+\frac{\widetilde{\sigma}_{2}^{2}}{1-\rho},$ (3.25)
$\sigma_{Pj}^{2}=$ $\textsf{E}\Big(\textsf{Pj}\big[\frac{Y(1)-\mu_{1}}{\rho}+\frac{Y(2)-\mu_{2}}{1-\rho}\big|\bm{\phi}(\bm{X})\big]\Big)^{2},$
$\widetilde{\sigma}_{t}^{2}=$ $\textsf{E}\Big(Y(t)-\mu_{t}-\textsf{Pj}\big[Y(t)-\mu_{t}\big|\bm{\phi}(\bm{X})\big]\Big)^{2}$
$+\rho(1-\rho)\textsf{E}\Big(\textsf{Pj}\big[Y(1)-\mu_{1}-(Y(2)-\mu_{2})\big|\bm{\phi}(\bm{X})\big]\Big)^{2},\;t=1,2.$
@@
The test always controls the type I error rate but is conservative when $\sigma_{\mathcal{T}}<1$ ($\textsf{Pj}\big[(Y(1)-\mu_{1})/\rho+(Y(2)-\mu_{2})/(1-\rho)\big|\bm{\phi}(\bm{X})\big)\not\equiv 0$). Under the local alternatives, the asymptotic power is
@@
If we replace $\widehat{\sigma}_{n,t}^{2}$ in (3.22) by the consistent estimators $\widehat{\widetilde{\sigma}}_{n,t}^{2}$ of $\widetilde{\sigma}_{t}^{2}$, $t=1,2$, and define the adjusted test statistic as
$\mathcal{T}_{adj}(n)=\frac{\overline{Y}_{n,1}-\overline{Y}_{n,2}}{\sqrt{\widehat{\widetilde{\sigma}}_{n,1}^{2}/N_{n,1}+\widehat{\widetilde{\sigma}}_{n,2}^{2}/N_{n,2}}}.$ (3.28)
@@
When $T_{i},\bm{\phi}(\bm{X}_{i})$, $i=1,\ldots,n$, are available in the analysis stage. We can obtain the consistent estimators as follow. Let $T_{i,1}=T_{i}$, $T_{i,2}=1-T_{i}$. We regress the residuals $T_{i,t}(Y_{i}-\overline{Y}_{n,1})$s over $T_{i,t}\bm{\phi}(\bm{X}_{i})$s by fitting the linear model
@@
$\widehat{\widetilde{\sigma}}_{n,t}^{2}=\frac{1}{\sum_{i=1}^{n}T_{i}(t)-1}\sum_{i=1}^{n}\big(\widehat{\zeta}_{i,t}\big)^{2}+\frac{\rho(1-\rho)}{n-2}\sum_{i=1}^{n}\big(\bm{\phi}(\bm{X}_{i})[\widehat{\bm{\alpha}}_{n,1}-\widehat{\bm{\alpha}}_{n,2}]^{\prime}\big)^{2},$ (3.29)
$t=1,2$.
@@
$\widehat{\widetilde{\sigma}}_{n,t}^{2}\overset{P}{\to}\widetilde{\sigma}_{t}^{2},\;t=1,2.$
3.1 and 3.2 will follow from (2.7) and (2.8). When the stratified randomization is adopted, Theorems 3.1 and 3.2 remain true even when $\gamma=0$ by Theorem 2.4.
It is seen that the asymptotic variance $\vec{\sigma}_{Z}^{2}$ in (2.8) has the closed form $\textsf{E}(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})])^{2}$ and does not exceed $\textsf{E}Z^{2}$ is important because in this case the classical test is valid and the variance is estimable so that the test can be adjusted to have the precise type I error rate.
When the stratified randomization is adopted, we always have $\vec{\sigma}_{Z}^{2}=\textsf{E}(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})])^{2}$. This is the main reason why the classical test under the stratified randomization is always valid and can be adjusted in varies ways.
@@ 4.1 The properties of the CAR procedures
We write $\bm{\phi}_{n}=\bm{\phi}(\bm{X}_{n})$, $\bm{\phi}=\bm{\phi}(\bm{X})$. Let $\Sigma=\textsf{E}\left[\bm{\phi}^{\prime}\bm{\phi}\right]$. Without loss of generality, we assume $\Sigma$ is non-singular. Otherwise, there is an orthogonal matrix $U$ such that $\Sigma=U^{\prime}diag(\sigma_{1}^{2},\ldots,\sigma_{q_{1}}^{2},0,\ldots)U$, where $\sigma_{t}^{2}$s are positive numbers. Let $\widetilde{\bm{\phi}}=\bm{\phi}U^{\prime}$. Then $\textsf{E}[\widetilde{\bm{\phi}}^{\prime}\widetilde{\bm{\phi}}]=diag(\sigma_{1}^{2},\ldots,\sigma_{q_{1}}^{2},0,\ldots)$, $\bm{\Lambda}_{n}=\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{\bm{\phi}}_{i}U$, $\|\bm{\Lambda}_{n}\|^{2}=\|\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{\bm{\phi}}_{i}\|^{2}$, $\langle\bm{\Lambda}_{n},\bm{\phi}_{n-1}\rangle=\langle\sum_{i=1}^{n}(T_{i}-\rho)\widetilde{\bm{\phi}}_{i},\widetilde{\bm{\phi}}_{n+1}\rangle$. So, it is sufficient to consider $\widetilde{\bm{\phi}}=(\widetilde{\phi}^{(1)},\ldots,\widetilde{\phi}^{(q)})$. Since $\textsf{P}(\widetilde{\phi}^{(t)}=0,t=q_{1}+1,\ldots q)=1$, it is sufficient to consider the subverter $\widetilde{\bm{\psi}}=(\widetilde{\phi}^{(1)},\ldots,\widetilde{\phi}^{(q_{1})})$ with $\textsf{E}[\widetilde{\bm{\psi}}^{\prime}\widetilde{\bm{\psi}}]=diag(\sigma_{1}^{2},\ldots,\sigma_{q_{1}}^{2})$ being non-singular.
Let $\mathscr{F}_{n}=\sigma(\bm{X}_{i},T_{i},Z_{i},W_{i},i=1,2,\ldots,n)$ be the history sigma generated by results of variables and allocations up to stage, $\ell_{n}=\ell\left(\frac{\langle\bm{\Lambda}_{n-1},\bm{\Phi}_{n}\rangle}{(n-1)^{\gamma}}\right)$. Then, $\textsf{P}(T_{n+1}=1|\mathscr{F}_{n},\bm{X}_{n+1})=\ell_{n+1}$ by (2.1). Let $r\in[2,r_{0}]$. Write $\bm{\psi}_{n}=(T_{i}-\rho)\bm{\phi}(\bm{X}_{n})$. Then $\bm{\Lambda}_{n+1}=\bm{\Lambda}_{n}+\bm{\psi}_{n}$. From the elementary inequality that

### v3

@@ 2.2 Theoretical Properties
Suppose $\textsf{E}Z^{2}<\infty$, $\textsf{E}W^{2}<\infty$, $\textsf{E}W=0$, and let $\ell(x)$ be defined as in (2.13). Then
@@
for all $r>0$. Moreover, for all the case of $0<\gamma<1$, (2.11) and (2.12) hold.
@@
Let $D_{n}=\sum_{i=1}^{n}(T_{i}-\rho)$, $D_{n}(l,k_{l}^{\ast})=\sum_{i=1}^{n}(T_{i}-\rho)I\{X_{i,l}=x_{l}^{k_{l}^{\ast}}\}$, $D_{n}(k_{1}^{\ast},\ldots,k_{p}^{\ast})=\sum_{i=1}^{n}(T_{i}-\rho)I\{X_{i,1}=x_{1}^{k_{1}^{\ast}},\ldots,X_{i,p}=x_{p}^{k_{p}^{\ast}}\}$. Then
@@
When $w_{s}\neq 0$, the procedure is related to the stratified randomization. When $w_{s}=0$, $w_{m,l}\neq 0$, $l=1,\ldots,p$, the procedure is related to the Pocock and Simon, (1975)’s marginal procedure. When $w_{s}=w_{m,l}=0$, $l=1,\ldots,p$, the procedure is related to Efron, (1971)’s biased coin design.
@@
Thus, when $w_{s}\neq 0$, $D_{n}(k_{1},\ldots,k_{p})=O_{P}(n^{\gamma/2})$ and so, for any $m(\bm{X})$, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})=O_{P}(n^{\gamma/2})$. When $w_{s}=0$, we will have
$D_{n}(l,k_{l})=O_{P}(n^{\gamma/2})\;\text{ and }$ $\sum_{i=1}^{n}(T_{i}-\rho)f(X_{i,l})=O_{P}(n^{\gamma/2}),\text{ if }w_{m,l}\neq 0,$
@@
with $\vec{\sigma}_{m}^{2}=\vec{\sigma}_{m-\textsf{E}m}^{2}+\vec{\sigma}_{\textsf{E}m}^{2}=\vec{\sigma}_{m-\textsf{E}m}^{2}\leq\rho(1-\rho)\textsf{Var}(m(\bm{X}))$ since $1\in Span(\bm{\phi}(\bm{X}))$. In particular,
@@
When $\rho=1/2$, the allocation probability is similar to that of Zhang, (2021), where (2.17) is obtained, but the properties (2.18) and (2.19) are not found.
When $\rho=1/2$ and $\gamma=0$, our CAR procedure is similar to that of Hu and Hu, (2012). The theory of this kind of CAR procedures can be found in Hu and Zhang, (2020) and Zhang, (2023). It is shown that (2.18) still holds, but $\vec{\sigma}_{m}^{2}$ has no closed from and may exceed $\textsf{E}[m^{2}(\bm{X})]/4$ unless $\textsf{E}[m(\bm{X})|\bm{\phi}(\bm{X})]\in Span\{\bm{\phi}(\bm{X})\}$. In particular, when $w_{s}=w_{0}=0$, the procedure is Pocock and Simon, (1975)’s marginal procedure. Zhao et al., (2024) demonstrated via examples that along some directions, there is an increased variance of within-stratum imbalances under Pocock and Simon, (1975)’s procedure compared to the simple randomization.
@@
(2.5) and (2.6) hold with $\gamma=0$, i.e. $\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\right)^{1/r}\leq C_{r,r_{0}}n^{\frac{1}{r(r_{0}+1-r)}}$, $r\in[1,r_{0}]$. In particular, $\textsf{E}[\|\bm{\Lambda}_{n}\|^{2}]=O\big(n^{\frac{1}{r_{0}-1}}\big)$, and $O\big(n^{\frac{8}{(r_{0}+1)^{2}}}\big)$ when $r_{0}\geq 3$.
@@
If $\textsf{E}Z^{2}<\infty$, $\textsf{E}[Z|\bm{\phi}(\bm{X})]\in Span\{\bm{\phi}(\bm{X})\}$, and $\big(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})]\big)^{2}\in Span\{\bm{\phi}(\bm{X})\}$ when $\rho\neq 1/2$, then
@@ 3 An application
As an application, we study hypothesis testing of treatment effects under the proposed covariate-adaptive procedure. For each unit $i$, denote $Y_{i}(1)$ and $Y_{i}(2)$ as the potential outcomes under the treatments $1$ and $2$, respectively, with $\textsf{E}[Y_{i}(t)]=\mu_{t}$ and $\textsf{Var}(Y_{i}(t))=\sigma_{t}^{2}$, $t=1,2$. The observed outcome is $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(2).$ The treatment effect is defined as $\tau=\textsf{E}\{Y_{i}(1)\}-\textsf{E}\{Y_{i}(2)\}$. Also, denote $\bm{X}_{i}=(X_{i,1},\ldots,X_{i,p})$ as the covariate of unit $i$ as before. Assume that $(Y_{i}(1),Y_{i}(2),\bm{X}_{i})$, $i=1,\ldots,n$, are independent and identically distributed as $(Y(1),Y(2),\bm{X})$. We suppose that the units are randomized sequentially by the proposed CAR procedure with feature $\bm{\phi}(\bm{X})$. We consider the test of
@@
Since $\overline{Y}_{n,1}=\frac{1}{N_{n,1}}\sum_{i=1}^{n}T_{i}Y_{i}$, $\overline{Y}_{n,2}=\frac{1}{N_{n,2}}\sum_{i=1}^{n}(1-T_{i})Y_{i}$, $\widehat{\sigma}_{n,1}^{2}=\frac{1}{N_{n,1}-1}\sum_{i=1}^{n}T_{i}(Y_{i}-\overline{Y}_{n,1})^{2}$, $\widehat{\sigma}_{n,2}^{2}=\frac{1}{N_{n,2}-1}\sum_{i=1}^{n}(1-T_{i})(Y_{i}-\overline{Y}_{n,2})^{2}$ are consistent estimators of $\mu_{1}$, $\mu_{2}$, $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$, respectively, where $N_{n,1}=\sum_{i=1}^{n}T_{i}$ and $N_{n,2}=n-N_{n,1}$, we consider the classical test statistic
@@
Suppose that the proposed CAR procedure with a feature $\bm{\phi}(\bm{X})$ is applied to balance covariates between treatments $1$ and $2$ with a ratio $\rho:(1-\rho)$, and Assumptions 2.1, 2.2 for all $r_{0}>2$ are satisfied. Then
@@
$\sigma_{\tau}^{2}=\frac{\sigma_{1}^{2}}{\rho}+\frac{\sigma_{2}^{2}}{1-\rho}-\rho(1-\rho)\sigma_{Pj}^{2}=\frac{\breve{\sigma}_{1}^{2}}{\rho}+\frac{\breve{\sigma}_{2}^{2}}{1-\rho}+\breve{\sigma}_{3}^{2},$ (3.26)
$\sigma_{Pj}^{2}=$ $\textsf{E}\Big(\textsf{Pj}\big[\frac{Y(1)-\mu_{1}}{\rho}+\frac{Y(2)-\mu_{2}}{1-\rho}\big|\bm{\phi}(\bm{X})\big]\Big)^{2},$ (3.27)
$\breve{\sigma}_{t}^{2}=$ $\textsf{E}\Big(Y(t)-\mu_{t}-\textsf{Pj}\big[Y(t)-\mu_{t}\big|\bm{\phi}(\bm{X})\big]\Big)^{2},\;\;t=1,2,$
$\breve{\sigma}_{3}^{2}=$ $\textsf{E}\Big(\textsf{Pj}\big[Y(1)-\mu_{1}-(Y(2)-\mu_{2})\big|\bm{\phi}(\bm{X})\big]\Big)^{2}.$
@@
The test always controls the type I error rate but is conservative when $\sigma_{\mathcal{T}}<1$ ($\textsf{Pj}\big[(Y(1)-\mu_{1})/\rho+(Y(2)-\mu_{2})/(1-\rho)\big|\bm{\phi}(\bm{X})\big]\not\equiv 0$). Under the local alternatives, the asymptotic power is
@@
If $\widehat{\breve{\sigma}}_{n,t}^{2}$ are the consistent estimators of $\breve{\sigma}_{t}^{2}$, $t=1,2,3$, and we define the adjusted test statistic as
$\mathcal{T}_{adj}(n)=\frac{\overline{Y}_{n,1}-\overline{Y}_{n,2}}{\sqrt{\widehat{\breve{\sigma}}_{n,1}^{2}/N_{n,1}+\widehat{\breve{\sigma}}_{n,2}^{2}/N_{n,2}+\widehat{\breve{\sigma}}_{n,3}^{2}/n}}.$ (3.30)
@@
When $T_{i},\bm{\phi}(\bm{X}_{i})$, $i=1,\ldots,n$, are available in the analysis stage. We can obtain the consistent estimators as follow. Let $T_{i,1}=T_{i}$, $T_{i,2}=1-T_{i}$. We regress the residuals $T_{i,t}(Y_{i}-\overline{Y}_{n,t})$s over $T_{i,t}\bm{\phi}(\bm{X}_{i})$s by fitting the linear model
@@
$\widehat{\breve{\sigma}}_{n,t}^{2}=\frac{1}{\sum_{i=1}^{n}T_{i}(t)-1}\sum_{i=1}^{n}\big(\widehat{\zeta}_{i,t}\big)^{2},\;t=1,2,$ (3.31)
$\widehat{\breve{\sigma}}_{n,3}^{2}=\frac{1}{n-2}\sum_{i=1}^{n}\big(\bm{\phi}(\bm{X}_{i})[\widehat{\bm{\alpha}}_{n,1}-\widehat{\bm{\alpha}}_{n,2}]^{\prime}\big)^{2}.$
@@
$\widehat{\breve{\sigma}}_{n,t}^{2}\overset{P}{\to}\breve{\sigma}_{t}^{2},\;t=1,2,3.$
Theorems 3.1 and 3.2 will follow from (2.9) and (2.11).
It is important that the asymptotic variance $\vec{\sigma}_{Z}^{2}$ in (2.11) has the closed form $\textsf{E}(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})])^{2}$ and does not exceed $\textsf{E}Z^{2}$ because in this case the classical test is valid and the variance is estimable so that the test can be adjusted to have the precise type I error rate.
When the stratified randomization is adopted, Theorems 3.1 and 3.2 remain true even when $\gamma=0$ by Theorem 2.4. Under the stratified randomization, we always have $\vec{\sigma}_{Z}^{2}=\textsf{E}(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})])^{2}$. This is the main reason why the classical test under the stratified randomization is always valid and can be adjusted in varies ways.
@@ 4.1 The properties of the CAR procedures
We write $\bm{\phi}_{n}=\bm{\phi}(\bm{X}_{n})$, $\bm{\phi}=\bm{\phi}(\bm{X})$. Let $\Sigma=\textsf{E}\left[\bm{\phi}^{\prime}\bm{\phi}\right]$. Without loss of generality, we assume that $\Sigma$ is non-singular. In fact, if $\Sigma=\bm{0}$, then $\bm{\phi}=\bm{\phi}_{n}=\bm{\Lambda}_{n}=0$ a.s., and $\ell_{n}\equiv\rho$, (2.8)-(2.12), (2.14)-(2.15) are obvious. If $\Sigma\neq\bm{0}$, then there is an orthogonal matrix $U$ such that $\Sigma=Udiag(\sigma_{1}^{2},\ldots,\sigma_{q_{\ast}}^{2},0,\ldots)U^{\prime}$, where $\sigma_{t}^{2}$s are positive numbers. Let ${\bm{\psi}}=\bm{\phi}U^{\prime}$ and ${\bm{\psi}}_{i}=\bm{\phi}_{i}U^{\prime}$. Then $\textsf{E}[{\bm{\psi}}^{\prime}{\bm{\psi}}]=diag(\sigma_{1}^{2},\ldots,\sigma_{q_{\ast}}^{2},0,\ldots)$, $\bm{\Lambda}_{n}=\sum_{i=1}^{n}(T_{i}-\rho){\bm{\psi}}_{i}U$, $\|\bm{\Lambda}_{n}\|^{2}=\|\sum_{i=1}^{n}(T_{i}-\rho){\bm{\psi}}_{i}\|^{2}$, $\langle\bm{\Lambda}_{n},\bm{\phi}_{n+1}\rangle=\langle\sum_{i=1}^{n}(T_{i}-\rho){\bm{\psi}}_{i},{\bm{\psi}}_{n+1}\rangle$. So, it is sufficient to consider ${\bm{\psi}}=({\psi}^{(1)},\ldots,{\psi}^{(q)})$. Since $\textsf{P}({\psi}^{(t)}=0,t=q_{\ast}+1,\ldots q)=1$, it is sufficient to consider the subverter ${\bm{\psi}}_{\ast}=({\psi}^{(1)},\ldots,{\psi}^{(q_{\ast})})$ with $\textsf{E}[{\bm{\psi}}_{\ast}^{\prime}{\bm{\psi}}_{\ast}]=diag(\sigma_{1}^{2},\ldots,\sigma_{q_{\ast}}^{2})$ being non-singular.
Let $\mathscr{F}_{n}=\sigma(\bm{X}_{i},T_{i},Z_{i},W_{i},i=1,2,\ldots,n)$ be the history sigma field generated by results of variables and allocations up to stage $n$. Then, $\textsf{P}(T_{n+1}=1|\mathscr{F}_{n},\bm{X}_{n+1})=\ell_{n+1}$ by (2.1), where $\ell_{n}=\ell\left(\frac{\langle\bm{\Lambda}_{n-1},\bm{\Phi}_{n}\rangle}{(n-1)^{\gamma}}\right)$.
Let $r\in[2,r_{0}]$. Write $\bm{\psi}_{n}=(T_{n}-\rho)\bm{\phi}(\bm{X}_{n})$. Then $\bm{\Lambda}_{n+1}=\bm{\Lambda}_{n}+\bm{\psi}_{n}$. From the elementary inequality that